OpenStudy (a1234):

A bag holds 20 yellow mints and 80 other green or pink mints. You choose a mint at random, eat it, and choose another. Find the number of pink mints if P(yellow then pink) = P(green then yellow). What is the least number of pink mints if P(yellow then pink) = P(green then yellow)?

OpenStudy (anonymous):

sry i thought ik nvm

OpenStudy (kropot72):

Let G be the number of green mints and P be the number of pink mints. Then \[\frac{20}{100}\times\frac{P}{99}=\frac{G}{100}\times\frac{20}{99}\] from which we find that P = G. But G + P = 80, so 2P = 80 and \[p=\frac{80}{2}=40\ mints\]